OpenStudy (zaxoanl):

can anyone help me on integration ∫dx/(x(lnx)^5)

OpenStudy (anonymous):

Did you try a u - sub?

OpenStudy (anonymous):

u = lnx

OpenStudy (anonymous):

Yes, that should work :)\[\int\limits \frac{ dx }{ x(lnx)^5 }\]

OpenStudy (anonymous):

\[u = lnx, ~~~du = \frac{ 1 }{ x }dx\]

OpenStudy (anonymous):

Do you know how to do the rest?

OpenStudy (anonymous):